\begin{thebibliography}{99}
%
%\bibitem{[Aled15]} J.~A.~Aledo and R.~M.~Rubio, {\em On the scalar curvature of space-like hypersurfaces in %Generalized Robertson-Walker space-times}, Diff. Geom. and its applications , October 2015.
%
\bibitem{[Ali951]} 
L. Al\'{\i}as, A. Romero, and M. S\'anchez, {\em Uniqueness of complete spacelike hypersurfaces of constant mean curvature in generalized Robertson-Walker space-times}, Gen. Relativ. Gravit. {\bf 27} (1) (1995), 71--84.
%
%\bibitem{[Ali952]} 
%L. Al{\'\i}as, A. Romero, M. S\'anchez, Compact spacelike hypersurfaces of constant mean curvature in generalized %Robertson-Walker spacetimes. In: Dillen F. editor. Geometry and Topology of Submanifolds VII. River Edge NJ, %USA: World Scientific, 1995, pp 67--70.
%
%\bibitem{[Ali97]} 
%L.J. Alias, A. Romero, M. S‡nchez, {\em Spacelike hypersurfaces of constant mean curvature and Calabi-Bernstein %type problems}, Toh™ku Math. J. {\bf 49} (1997) 339--345.
%
%\bibitem{[Ali3]} 
%L. Al{\'\i}as, A. Romero, and M. S\'anchez, Space-like hypersurfaces of constant mean curvature in space-times %with symmetries
%
\bibitem{[Ars14]} 
K.~Arslan, R.~Deszcz, R.~Ezentas, M.~Hotlo\'s, C.~Murathan, {\em On generalized Robertson-Walker spacetimes satisfying some curvature condition}, Turk. J. Math. {\bf 38} (2014) 353--373.
%
\bibitem{[Ban14]}
B.-Y.~Chen, {\em A simple characterization of generalized Robertson-Walker space-times}, Gen. Relativ. Gravit. {\bf 46} (2014) 1833.
 %
%\bibitem{[Bar]} 
%A.~Barnes, B.~Edgar, and R.~Rani, {\em Killing Tensors from conformal Killing vectors}, arXiv:gr-qc/0212016v1, z%3Dec 2001.
%
%\bibitem{[Bee96]} 
%J.~K.~Beem, P.~E.~Ehrlich, K.~L.~Easley, Global Lorentzian Geometry, 2nd ed. Pure and Applied mathematics, %vol. 202, 1996 Marcel Dekker , New York.
%
%\bibitem{[BV]}
%M.~Brozos-V\'azquez, E.~Garcia-Rio, and R.~V\'azquez-Lorenzo, {\em Some remarks on locally conformally flat %static space-times}, J. Math. Phys. {\bf 46} (2005) 022501.
%
%\bibitem{[Cab]} 
%M.~Caballero, A.~Romero and R.~Rubio, {\em Constant mean curvature space-like hypersurfaces in Lorentzian %manifolds with a time-like gradient conformal vector field}, Class. Quantum Grav. {\bf 28} n.14 (2011) 145009 (14 pp.)
%
\bibitem{[Fia]} 
A.~Fialkow, {\em Conformal geodesics}, Trans. Amer. Math. Soc. {\bf 45} n.3 (1939) 443--473.
%
\bibitem{[Gut]} 
M. Guti\'errez and B. Olea, {\em Global decomposition of a Lorentzian manifold as a generalized Robertson-
Walker space}, Differ. Geom. Applic. {\bf 27} (2009) 146--156.
%
\bibitem{Hui13}
S.~K.~Hui, {\em On $\Phi$-pseudosymmetries of $(LCS)_n$-manifolds}, Kyungpook Math. J. {\bf 53} (2013)
285--294.
%
\bibitem{[Sur]}
C.~A.~Mantica and L.~G.~Molinari, {\em Generalized Robertson-Walker space-time, a survey}, Int. J. Geom. Meth. Mod. Phys. {\bf 14} (2017) 1730001 (27 pp.).
%
\bibitem{MM16} 
C.~A.~Mantica and L.~G.~Molinari, {\em On the Weyl and the Ricci tensors of Generalized Robertson-Walker 
space-times}, J. Math. Phys. {\bf 57} n.10 (2016), 102502.
%
%\bibitem{[MSD]} 
%C.~A.~Mantica , Y.~J.~Suh and U.~C.~De, {\em A note on generalized Robertson-Walker space-times}, 
%Int. J. Geom. Meth. Mod. Phys. 13 (2016) (9 pp).
%
\bibitem{[Rom1]} 
A. Romero, R.N. Rubio, and J.J. Salamanca, {\em Uniqueness of complete maximal hypersurfaces in spatially parabolic generalized Robertson-Walker space-times}, Class. Quantum Grav. {\bf 30} (11) (2013) 115007.
%
%\bibitem{[Rom2]} 
%A. Romero, R.N. Rubio, and J.J. Salamanca, {\em Uniqueness of complete maximal hypersurfaces in spatially %parabolic generalized Robertson-Walker space-times. Applications to uniqueness results}, Int. J. Geom. Meth. %Mod. Phys. 10 n.8 (2013),1360014.
%
\bibitem{[San98]} 
M. S\'anchez, {\em On the geometry of generalized Robertson-Walker space-times: geodesics}, Gen. Relativ. Gravit. {\bf 30} (1998), 915--932.
%
\bibitem{[San99]} 
M.~S\'anchez, {\em On the geometry of generalized Robertson-Walker space-times: curvature and Killing fields}, Gen. Relativ. Gravit. {\bf 31} (1999),1-15.
%
\bibitem{Shaikh03}
A.~A.~Shaikh, {\em On Lorentzian almost paracontact manifolds with a structure of the concircular type}, 
Kyungpook Math. J. {\bf 43} (2003) 305--314.
%
\bibitem{ShBa05}
A.~A.~Shaikh and K.~K.~Baishya, {\em On concircular structure spacetimes}, J. Math. Stat. {\bf 1} (2005) 129--132.
%
\bibitem{ShBa06}
A.~A.~Shaikh and K.~K.~Baishya, {\em On concircular structure spacetimes II}, American J. Appl. Sci. {\bf 3} n.4 (2006) 1790--1794.
%
\bibitem{Sh09}
A.~A.~Shaikh, {\em Some results on (LCS)$_n$-manifolds}, J. Korean Math. Soc. {\bf 46} (2009), 449--461.
%
\bibitem{[Yan1]} 
K. Yano, {\em Concircular geometry I-IV}, Proc. Imp. Acad. Tokyo {\bf 16} (1940) 195--200, 354--360, 442--448, 505--511.
%
\bibitem{[Yan2]} 
K.~Yano, {\em On the torseforming direction in Riemannian Spaces}, Proc. Imp. Acad. Tokyo {\bf 20} (1944) 340--345.
\end{thebibliography}

\newcommand{\rem}[1]{}
\documentclass{amsart}
\usepackage{amsfonts,amssymb,amsmath,amsthm,mathrsfs}
\usepackage{url}
\usepackage[dvips]{epsfig}
\newtheorem{thrm}{Theorem} 

\newtheorem{prop}[thrm]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[thrm]{Definition}
\begin{document}
\author[C.~A.~Mantica and L.~G.~Molinari]
{Carlo alberto Mantica and Luca~Guido~Molinari}
\address{L.~G.~Molinari (corresponding author): Physics Department,
Universit\`a degli Studi di Milano and I.N.F.N. sez. Milano,
Via Celoria 16, 20133 Milano, Italy -- C.~A.~Mantica: I.I.S. Lagrange, Via L. Modignani 65, 
20161, Milano, Italy.}
\email{luca.molinari@unimi.it, carloalberto.mantica@libero.it}
\subjclass[2010]{Primary 53B30, Secondary 53C25}
\keywords{Generalized Robertson-Walker space-time, Lorentzian concircular structure,
torse-forming vector, concircular vector.}
\title[]{A note on Concircular Structure space-times}
\begin{abstract}
In this note we show that Lorentzian Concircular Structure manifolds $(LCS)_n$ coincide with Generalized Robertson-Walker space-times.
\end{abstract}
\date{1 april 2018}
\maketitle
%
%
%

{\sf Generalized Robertson-Walker} (GRW) space-times were introduced in 1995 by Al{\'\i}as, Romero and S\'anchez \cite{[Ali951]} as the warped product $-1\times_{q^2} M^*$, 
where $(M^*,g^*)$ is a Riemannian submanifold. In other terms, they are Lorentzian manifolds 
characterised by a metric 
\begin{align}\label{(1.1)}
g_{ij}dx^i dx^j = -(dt)^2 + q(t)^2 g^*_{\mu\nu}(x^1,...,x^{n-1}) dx^\mu dx^\nu
\end{align}
They are interesting not only for geometry \cite{[San98],[San99],[Gut],[Rom1],[Ars14]}, but also for physics: they include relevant space-times such as Robertson-Walker, Einstein-de Sitter, static Einstein, de-Sitter, 
the Friedmann cosmological models. They are a wide generalization of space-times for cosmological models.

In 2003 A. A. Shaikh \cite{Shaikh03} introduced the notion of {\sf Lorentzian Concircular Structure} $(LCS)_n$. It is a Lorentzian manifold endowed with a unit time-like concircular vector field, i.e. $u^i u_i=-1$ and
\begin{align}\label{(1.2)}
\nabla_k u_j = \varphi (u_ku_j+g_{kj}) 
\end{align}
where $\varphi\neq 0$ is scalar function obeying
\begin{align}\label{(1.3)}
\nabla_j \varphi =\mu u_j
\end{align}
being $\mu $ a scalar function. Various authors studied the properties of $(LCS)_n$ manifolds \cite{ShBa05,ShBa06,Sh09,Hui13}.

We show that GRW and $(LCS)_n$ are the same space-times.\\

We recall few definitions that will be used in this note. The first ones are the definitions of ``torse-forming'' and
``concircular'' vector fields, by Yano:
\begin{definition}[Yano, \cite{[Yan1],[Yan2]}]
A vector field $X_j$ is named torse-forming if $\nabla_k X_j = \omega_k X_j+\varphi g_{kj}$, being $\varphi $ 
a scalar function and $\omega_k$ a non vanishing one-form.
It is named concircular if $\omega_k$ is a gradient or locally a gradient of a scalar function.
\end{definition}
\noindent
Fialkow gave a definition different from Yano's: 
\begin{definition}[Fialkow \cite{[Fia]}]
A vector field $X_j$ is named concircular if it satisfies $\nabla_k X_j =  \rho g_{kj}$,
being $\rho $ a scalar function.
\end{definition}

The following simple but deep result was recently proven:
\begin{thrm}[Bang-Yen Chen, \cite{[Ban14]}]
A $n>3$ dimensional Lorentzian manifold is a GRW space-time if and only if it admits a time-like concircular vector 
field (in the sense of Fialkow). 
\end{thrm}
It is worth noticing that for a unit time-like vector field, the torse-forming property by Yano becomes precisely
eq.\eqref{(1.2)}, with generic scalar field $\varphi $.
%
Based on Chen's theorem we proved:
\begin{prop}[Mantica and Molinari, \cite{[Sur],MM16}]
A $n>3$ dimensional Lorentzian manifold is a GRW space-time if and only if it admits a unit time-like 
torse-forming vector, \eqref{(1.2)}, that is also an eigenvector of the Ricci tensor.
\end{prop}
Now comes the equivalence: from \eqref{(1.2)} (holding either for GRW and $(LCS)_n$ space-times)
we evaluate 
$$R_{jkl}{}^m u_m = [\nabla_j,\nabla_k]u_l = (h_{kl} \nabla_j - h_{jl}\nabla_k) \varphi -\varphi^2
(u_j g_{kl}-u_k g_{jl}) $$
where $h_{kl}= u_ku_l + g_{kl}$. Contraction with $g^{jl}$ gives
\begin{align}
 R_k{}^m u_m = u_k [u^m\nabla_m \varphi +(n-1)\varphi^2 ] - (n-2) \nabla_k \varphi 
 \end{align}
If \eqref{(1.3)} holds, then $u_k$ is an eigenvector of the Ricci tensor, and we conclude that {\em a $(LCS)_n$ manifold is a GRW space-time}.\\
If $R_{km} u^m =\xi u_k$ it is $(n-2) \nabla_k \varphi = \alpha u_k$ for some scalar field $\alpha $, i.e. \eqref{(1.3)}
holds. Then we conclude that {\em a GRW space-time is a $(LCS)_n$ manifold}. 
\vfill
\end{document}